# RIDE: Difficulty Evolving Perturbation with Item Response Theory for Mathematical Reasoning

### v1

@@ Abstract
Large language models (LLMs) achieve high performance on mathematical reasoning, but these results can be inflated by training data leakage or superficial pattern matching rather than genuine reasoning. To this end, an adversarial perturbation–based evaluation is needed to measure true mathematical reasoning ability. Current rule-based perturbation methods often generate ill-posed questions and impede the systematic evaluation of question difficulty and the evolution of benchmarks. To bridge this gap, we propose RIDE, a novel adversarial question-rewriting framework that leverages Item Response Theory (IRT) to rigorously measure question difficulty and to generate intrinsically more challenging, well-posed variations of mathematical problems. We employ $35$ LLMs to simulate students and build a difficulty ranker from their responses. This ranker provides a reward signal during reinforcement learning and guides a question-rewriting model to reformulate existing questions across difficulty levels. Applying RIDE to competition-level mathematical benchmarks yields perturbed versions that degrade advanced LLM performance, with experiments showing an average $21.73\%$ drop across $26$ models, thereby exposing limited robustness in mathematical reasoning and confirming the validity of our evaluation approach. Code and data of our paper are available at https://github.com/LiXinyuan1015/RIDE.
@@ 1 Introduction
Mathematical reasoning is widely regarded as a crucial testament for evaluating the reasoning capabilities of large language models (LLMs). Recent works like DeepSeek-Math (shao2024deepseekmathpushinglimitsmathematical) and Qwen2.5-Math (yang2024qwen25mathtechnicalreportmathematical) clearly demonstrate the benefits of reinforcement learning for mathematical reasoning. Nevertheless, strong performance in mathematical reasoning may not directly reflect the model’s reasoning ability (wu2025reasoningmemorizationunreliableresults). It may instead result from data contamination during pretraining (golchin2024timetravelllmstracing) that leads to memorization-based answers (zhang2024carefulexaminationlargelanguage; li2025diagnosingmemorizationchainofthoughtreasoning), or from the model relying primarily on superficial pattern matching without truly understanding mathematical knowledge (li2024gsmpluscomprehensivebenchmarkevaluating; mirzadeh2025gsmsymbolicunderstandinglimitationsmathematical). This phenomenon undermines the reliability of LLMs as genuine mathematical reasoners.
@@
To address this issue, there is an urgent need for a more adversarial evaluation benchmark that can faithfully characterize the robustness of mathematical reasoning. For example, a rule-based rewriting method is proposed (li2024gsmpluscomprehensivebenchmarkevaluating; zhou2024modelreallygoodmath), which applies operations such as numerical substitution and symbol modification to reformulate questions. This method has been shown to reduce model performance. However, rule-based approaches may render the rewritten questions ill-posed and unsolvable (xue2025reliablemathbenchmarkreliablemathematical), thereby undermining their validity as reliable benchmarks. Moreover, previous work has focused mainly on low-difficulty datasets cobbe2021trainingverifierssolvemath, with limited attempts at more challenging data. Based on this, we identify the following challenges: (1) Robust Adversarial Perturbation. Robust perturbations to the data need to both degrade the model performance and remain credible and reasonable, which requires shifting from injecting noise to directly increasing the difficulty of the questions as a means of perturbation. Moreover, the perturbation should be unpredictable for changing, and simple rules cannot capture the evolution. (2) More Challenging Data. The perturbation of data should not be limited to the same difficulty level of mathematical reasoning. It is also necessary to explore its effectiveness on difficulty evolution, ensuring reasonable perturbations can be applied to mathematical competition questions. These challenges require us to perturb existing benchmarks in ways that substantively increase the intrinsic difficulty of the questions (huang2025mathperturbbenchmarkingllmsmath). A question’s difficulty is often defined by whether a model answers correctly (lan2024surveynaturallanguageprocessing), but this definition depends heavily on the selected models, and major voting across these models ignores the individual differences in capabilities. Besides, defining difficulty by the steps of the reasoning chain (yu-etal-2025-generating) is easily misled by shortcut reasoning and does not account for the numerical computation complexity specific to mathematical reasoning.
Figure 1: Answers of Qwen3-30B-A3B to a mathematical question and variation rewritten by RIDE.
@@
To this end, we propose RIDE, a benchmark perturbation rewriting framework that incorporates Item Response Theory (IRT) to measure question difficulty (vania2021comparingtestsetsitem; scarlatos2025smartsimulatedstudentsaligned). Specifically, $35$ cutting-edge LLMs are treated as students to answer $2,000$ questions sampled from the high-difficulty mathematical dataset DeepMath (he2025deepmath103klargescalechallengingdecontaminated). A statistical model jointly characterizes the latent abilities of the student models and the difficulty parameters of the questions. Based on the IRT-estimated difficulty, a pairwise ranker is trained: given question embedding vectors, it outputs the probability that one question is more difficult than another. The ranker’s output further serves as a difficulty reward signal for reinforcement learning on Qwen3-8B, resulting in the question rewriting model RIDE-8B. Applying reinforcement learning to difficulty-evolving question rewriting encourages the model to adaptively generate questions with higher difficulty, thus avoiding superficial modifications, while further enhancing consistency between the rewritten questions and their answers by setting a correctness reward. RIDE-8B is then applied to rewrite competition-level mathematical benchmarks such as AIME-24 and AMC-23, achieving difficulty evolution for these datasets. Experiments demonstrate that the perturbed benchmarks induce performance drops in frontier reasoning models, including GPT-5, DeepSeek-V3.1, and others, thereby enabling robust evaluation and comparison of advanced models’ mathematical reasoning abilities (as shown in Figure 1). In addition, RIDE can be used as a data augmentation method to enhance training data and improve mathematical reasoning performance. Our contributions are as follows:
@@
  • We propose a benchmark perturbation framework with IRT to evaluate robustness in mathematical reasoning, reducing partial data leakage during pretraining on evaluation. We measure robustness via performance degradation and observe that most LLMs are sensitive to perturbations, indicating poor robustness.
  • We propose a more scalable IRT-based difficulty estimation method and a corresponding difficulty ranker. In addition, we introduce a reinforcement learning scheme for training a question-rewriting model.
@@
  • We release a question rewriting model (RIDE-8B), two rewritten competition-level benchmarks (RIDE-AIME and RIDE-AMC), and a training dataset of over 10K rewritten questions (RIDE-DeepMath).
@@ 2 Related Work
Robustness Evaluation. Existing LLM robustness evaluations (ul-abedin-etal-2025-arithmattack) often use controlled perturbations to test whether models truly understand tasks rather than memorize surface patterns (li2024gsmpluscomprehensivebenchmarkevaluating; yang2025utmathmathevaluationunit). A common practice is to construct original items and perturbed counterparts around the same semantics and compare the resulting performance drop (goel2021robustnessgymunifyingnlp; chang2023drspiderdiagnosticevaluationbenchmark; hou2025automaticrobustnessstresstesting; yang-etal-2025-evaluating). For example, GSM-PLUS li2024gsmpluscomprehensivebenchmarkevaluating applies rule-based perturbations such as numerical substitution (shrestha2025mathematicalreasoninglargelanguage) and operation modification, revealing substantial degradation in many LLMs. Unlike rule-based methods, we adapt a reinforcement learning-based rewriting approach to generate perturbations, producing better-formed data, avoiding unsolvable or multiple-solution cases, and increasing unpredictability.
Item Response Theory (IRT). Originating from educational measurement, IRT is widely applied in AI for automated item difficulty estimation (inbook; Chen_2025) and efficient evaluation (polo2024tinybenchmarksevaluatingllmsfewer; Martínez-Plumed_Castellano_Monserrat-Aranda_Hernández-Orallo_2022). Traditional IRT depends on real student responses, which are often difficult to obtain due to privacy constraints. Prior work addresses this by simulating response matrices (lalor-etal-2019-learning; truong2025reliableefficientamortizedmodelbased) or predicting difficulty from text (Martínez-Plumed_Castellano_Monserrat-Aranda_Hernández-Orallo_2022), which leads to high computational costs (truong2025reliableefficientamortizedmodelbased). We mitigate these issues via data augmentation to reduce the number of required student models and reframing difficulty estimation as a pairwise ranking task to reduce numerical prediction errors.
Reinforcement Learning for Reasoning. Reinforcement learning (RL) can enhance LLMs’ reasoning by rewarding correct solution paths and penalizing errors (deepseekai2025deepseekr1incentivizingreasoningcapability). Methods like GRPO (shao2024deepseekmathpushinglimitsmathematical) and GSPO (zheng2025groupsequencepolicyoptimization) normalize rewards within groups of candidate solutions, yielding strong gains on reasoning ability. Building on this, we propose a reasoning-driven approach to developing a rewriting model with RL.
@@ 3 Preliminary
We briefly introduce the IRT-based difficulty in our method. Consider a test taker with fixed ability $\theta$ and a math reasoning question $q$ with a difficulty score $d$. The response $y$ is modeled as a Bernoulli random variable that indicates correctness, where $y=1$ denotes a correct answer and $y=0$ denotes an incorrect answer. We introduce the Rasch model (rasch1993probabilistic), a foundational one-parameter (1-PL) IRT model in which the probability of a correct response is the logistic function of the difference between the test taker’s ability $\theta$ and the item’s difficulty $d$:
@@
Given a binary response matrix $\mathbf{Y}\in\{0,1\}^{M\times N}$, where $M$ and $N$ denote the number of test takers and questions, respectively. Each entry $Y_{ij}$ indicates the $i$-th test taker’s answer to the $j$-th question. Based on this response matrix, one can estimate the ability parameters $\theta$ and the difficulty parameters $d$ using a variety of statistical inference methods, such as the Bayesian inference via Markov chain Monte Carlo (MCMC) sampling (chen2021itemresponsetheory) and variational inference (VI) (Blei_2017). In this paper, we apply the VI method provided by py-irt (Lalor_2023). Concretely, we fit the Rasch model via stochastic variational inference on the response matrix with a mean-field Normal guide for $\theta$ and $d$, maximizing the evidence lower bound $\mathcal{L}$:
@@
$\max_{g}\ \mathcal{L}=\sum_{i=1}^{M}\sum_{j=1}^{N}\mathbb{E}_{g}\!\left[\,Y_{ij}(\theta_{i}-d_{j})-\log\!\big(1+e^{\theta_{i}-d_{j}}\big)\right]-\mathrm{KL}\!\big(g\,\|\,p\big)$
@@ 4 Method
In this section, we elaborate on the details of RIDE-Math. Inspired by the IRT-based difficulty estimation method with LLMs acting as students (truong2025reliableefficientamortizedmodelbased), we construct a difficulty ranker for math questions by performing parameter estimation described in Section 3. This ranker serves as a reward model by providing difficulty reward signals, which are used to guide reinforcement learning on the existing high-difficulty dataset DeepMath (he2025deepmath103klargescalechallengingdecontaminated). Through this rewriting process, we are able to generate difficulty in evolving math reasoning data. The overall architecture of our method is shown in Figure 2
Figure 2: The overall architecture of RIDE.
@@ 4.1.1 IRT Parameter Estimation
  • Variational Autoencoder (VAE) method. We mitigate the scarcity of student responses by training a VAE (kingma2022autoencodingvariationalbayes) on the observed response matrix $\mathbf{Y}$ to learn a low-dimensional manifold of student behavior. During generation, we draw a latent code $\mathbf{z}$ from the Gaussian prior and decode it into per-question correctness probabilities $\pi(\mathbf{z})$, from which Bernoulli draws yield a synthetic student-question response vector with conditional independence across items given $\mathbf{z}$:
$\tilde{\mathbf{Y}}=\begin{bmatrix}\mathbf{Y}\\[2.0pt] \mathbf{R}_{\mathrm{VAE}}\\[2.0pt] \mathbf{R}_{\mathrm{Sampling}}\end{bmatrix},\quad\mathbf{R}_{\mathrm{VAE}}=[\mathbf{r}_{\mathrm{VAE}}^{1},...,\mathbf{r}_{\mathrm{VAE}}^{S}]^{\intercal},\quad\mathbf{R}_{\mathrm{Sampling}}=[\mathbf{r}_{\mathrm{Sampling}}^{1},...,\mathbf{r}_{\mathrm{Sampling}}^{T}]^{\intercal}$
$S$ and $T$ denote the number of synthetic students by VAE and sampling, respectively. Then the modeling ability of IRT calibration for question difficulties can be improved. We perform IRT parameter estimation based on the response matrix $\tilde{\mathbf{Y}}$. Following the parameter estimation method described in Section 3, we obtain an IRT-estimated difficulty $d$ for each question and construct $N$ question-difficulty pairs $(q,d)$ for training the difficulty ranker, where $q$ denotes the question.
@@ 4.1.2 Pairwise Difficulty Ranker
In traditional IRT, estimating the difficulty of a new question requires collecting fresh student response data. To endow IRT difficulty estimation with generalization to unseen questions, we adopt an amortized approach (truong2025reliableefficientamortizedmodelbased), which extracts content features via an embedding model and learn a mapping from question content to difficulty value. However, due to the domain-specific symbolic semantics of math questions (mirzadeh2025gsmsymbolicunderstandinglimitationsmathematical), directly regressing from embeddings to a scalar difficulty may result in unreliable performance. Therefore, instead of regressing the difficulty value, we frame the task as a pairwise ranking question, training a ranker to identify the more difficult question from a pair of question embeddings. We use OpenAI’s text-embedding-3-large  [^1] https://platform.openai.com/docs/models/text-embedding-3-large model to obtain embedding vectors for each question-difficulty pair $(q,d)$, creating $N$ question-embedding-difficulty triplets as $(q,e,d)$. We then group these pairs by fine-grained topics (gao2024omnimathuniversalolympiadlevel) like algebra, geometry, and so on. Within each subcategory, ranking data are generated by forming pairwise combinations of questions. Specifically, for any two questions $(q_{i},q_{j})$ within the same subcategory, we create the ordered pair $(e_{i},e_{j},r_{ij})$ and its symmetric counterpart $(e_{j},e_{i},1-r_{ij})$ to reduce positional bias, where $r_{ij}$ denotes to the ranking label: it equals to $1$ if the question $i$’s difficulty $d_{i}$ exceeds the question $j$’s difficulty $d_{j}$, and 0 otherwise. To account for the influence of the magnitude gap between difficulty scores $d$, we assign a weight to each ranking pair. For $(e_{j},e_{j},r_{ij})$, the weight is $w_{ij}=\sqrt{\lvert d_{i}-d_{j}\rvert}$. We use these pairwise samples to train a multilayer perceptron (MLP) as the difficulty ranker. The input $\mathbf{x}$ of MLP is a fusion of absolute and relative features, which is concatenated as:
@@
$\mathcal{L}_{\mathrm{BCE}}=\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\mathcal{B}}w_{ij}\,\Big(-r_{ij}\log p(i,j)-(1-r_{ij})\log\!\big(1-p(i,j)\big)\Big)$

### v2

@@ Abstract
Large language models (LLMs) achieve high performance on mathematical reasoning, but these can be inflated by training data leakage or superficial pattern matching rather than genuine reasoning. To this end, an adversarial perturbation–based evaluation is needed to measure true mathematical reasoning ability. Current rule-based perturbation methods often generate ill-posed questions and impede the systematic evaluation of question difficulty and the evolution of benchmarks. To bridge this gap, we propose RIDE, a novel adversarial question-rewriting framework that leverages Item Response Theory (IRT) to rigorously measure question difficulty and to generate intrinsically more challenging, well-posed variations of mathematical questions. We employ $35$ LLMs to simulate students and build a difficulty ranker from their responses. This ranker provides a reward signal during reinforcement learning and guides a question-rewriting model to reformulate existing questions across difficulty levels. Applying RIDE to competition-level mathematical benchmarks yields perturbed versions that degrade advanced LLM performance, with experiments showing an average $21.73\%$ drop across $26$ models, thereby exposing limited robustness in mathematical reasoning and confirming the validity of our evaluation approach.
Machine Learning, ICML
@@ 1 Introduction
Mathematical reasoning is regarded as a crucial testament for evaluating the reasoning capabilities of large language models (LLMs). Recent works like DeepSeek-Math (Shao et al., 2024) and Qwen2.5-Math (Yang et al., 2024) clearly demonstrate the benefits of reinforcement learning for mathematical reasoning. Nevertheless, strong performance in mathematical reasoning may not directly reflect the model’s reasoning ability (Wu et al., 2025). It may instead result from data contamination during pretraining (Golchin and Surdeanu, 2024) that leads to memorization-based answers (Zhang et al., 2024; Li et al., 2025), or from the model relying primarily on superficial pattern matching without understanding mathematical knowledge (Li et al., 2024; Mirzadeh et al., 2025). This phenomenon undermines the reliability of LLMs as genuine mathematical reasoners.
@@
To address this issue, there is an urgent need for a more adversarial evaluation benchmark that can faithfully characterize the robustness of mathematical reasoning. For example, a rule-based rewriting method is proposed (Li et al., 2024; Zhou et al., 2024), which applies operations such as numerical substitution and symbol modification to reformulate questions. This method has been shown to reduce model performance. However, rule-based approaches may render the rewritten questions ill-posed and unsolvable (Xue et al., 2025), thereby undermining their validity as reliable benchmarks. Moreover, previous work has focused mainly on low-difficulty datasets (Cobbe et al., 2021), with limited attempts at more challenging data. Based on this, we identify the following challenges: (1) Robust Adversarial Perturbation. Robust perturbations to the data need to both degrade the model performance and remain credible and reasonable, which requires shifting from injecting noise to reformulating questions into formats that are inherently more challenging for the model to resolve correctly. Moreover, the perturbation should be unpredictable for changing, and simple rules cannot capture the evolution. (2) More Challenging Data. The perturbation of data should not be limited to the same difficulty level of mathematical reasoning. It is also necessary to explore its effectiveness on difficulty evolution, ensuring reasonable perturbations can be applied to mathematical competition questions. These challenges require us to perturb existing benchmarks in ways that substantively increase the intrinsic difficulty of the questions (Huang et al., 2025). A question’s difficulty is often defined by whether a model answers correctly (Lan et al., 2024), but this definition depends heavily on the selected models, and majority voting across these models ignores the individual differences in capabilities. Besides, defining difficulty by the steps of the reasoning chain (Yu et al., 2025) is easily misled by shortcut reasoning and does not account for the numerical computation complexity specific to mathematical reasoning.
Figure 1: Difficulty evolution via RIDE. The original question (left) is successfully answered, whereas the rewritten question (right) with added constraints triggers a reasoning error of the Qwen3-30B-A3B model.
@@
To this end, we propose RIDE, a benchmark perturbation rewriting framework that incorporates Item Response Theory (IRT) to measure question difficulty (Vania et al., 2021; Scarlatos et al., 2025). Specifically, $35$ cutting-edge LLMs are treated as students to answer $2,000$ questions sampled from the high-difficulty mathematical dataset DeepMath (He et al., 2025). A statistical model jointly characterizes the latent abilities of the student models and the difficulty parameters of the questions. This difficulty is not human cognitive complexity but a model-centric metric representing the inherent challenge a specific question format poses to the model’s ability to respond correctly. Based on the IRT-estimated difficulty, a pairwise ranker is trained: given question embedding vectors, it outputs the probability that one question is more difficult than another. The ranker’s output further serves as a difficulty reward signal for reinforcement learning on Qwen3-8B, resulting in the question rewriting model RIDE-8B. Applying reinforcement learning to difficulty-evolving question rewriting encourages the model to adaptively generate questions with higher difficulty, thus avoiding superficial modifications, while further enhancing consistency between the rewritten questions and their answers by setting a correctness reward. RIDE-8B is then applied to rewrite competition-level mathematical benchmarks such as AIME-24 and AMC-23, achieving difficulty evolution for these datasets. Experiments demonstrate that the perturbed benchmarks induce performance drops in frontier reasoning models, including GPT-5, DeepSeek-V3.1, and others, thereby enabling robust evaluation and comparison of advanced models’ mathematical reasoning abilities (as shown in Figure 1). In addition, RIDE can be used as a data augmentation method to enhance training data and improve mathematical reasoning performance. Our contributions are as follows:
@@
  • We propose a benchmark perturbation framework with IRT to evaluate robustness in mathematical reasoning. We measure robustness via performance degradation and observe that most LLMs are sensitive to perturbations, indicating poor robustness.
@@
  • We propose an IRT-based difficulty ranker and introduce a reinforcement learning scheme for training a question-rewriting model. We also release our question rewriting model (RIDE-8B), two rewritten competition-level benchmarks (RIDE-AIME and RIDE-AMC), and a training dataset of over 10K rewritten questions (RIDE-DeepMath).
@@ 2 Related Work
Robustness Evaluation. Existing LLM robustness evaluations (Ul Abedin et al., 2025) often use controlled perturbations to test whether models truly understand tasks rather than memorize surface patterns (Li et al., 2024; Yang et al., 2025a). A common practice is to construct original items and perturbed counterparts around the same semantics and compare the resulting performance drop (Goel et al., 2021; Chang et al., 2023; Hou et al., 2025; Yang et al., 2025b; Sun et al., 2025). For example, GSM-PLUS (Li et al., 2024) applies rule-based perturbations such as numerical substitution (Shrestha et al., 2025) and operation modification, revealing substantial degradation in many LLMs. Unlike rule-based methods, we adapt a reinforcement learning-based rewriting approach to generate perturbations, producing better-formed data, avoiding unsolvable or multiple-solution cases, and increasing unpredictability.
Item Response Theory (IRT). Originating from educational measurement, IRT is widely applied in AI for automated item difficulty estimation (Feng et al., 2025; Chen and Shiu, 2025) and efficient evaluation (Polo et al., 2024; Martínez-Plumed_Castellano_Monserrat-Aranda_Hernández-Orallo_2022; Xu et al., 2025b). IRT depends on real student responses, which are often difficult to obtain due to privacy constraints. Prior work addresses this by simulating response matrices (Lalor et al., 2019; Truong et al., 2025) or predicting difficulty from text (Martínez-Plumed_Castellano_Monserrat-Aranda_Hernández-Orallo_2022), which leads to high computational costs (Truong et al., 2025). We mitigate these issues via data augmentation to reduce the number of required student models and reframing difficulty estimation as a pairwise ranking task to reduce numerical prediction errors.
@@ 3 Preliminary
We briefly introduce the IRT-based difficulty in our method. Consider a test taker with fixed ability $\theta$ and a math reasoning question $q$ with a difficulty score $d$. The response $y$ is modeled as a Bernoulli random variable that indicates correctness, where $y=1$ denotes a correct answer and $y=0$ denotes an incorrect answer. We introduce the Rasch model (Rasch, 1993), a foundational one-parameter (1-PL) IRT model in which the probability of a correct response is the logistic function of the difference between the test taker’s ability $\theta$ and the item’s difficulty $d$:
@@
Given a binary response matrix $\mathbf{Y}\in\{0,1\}^{M\times N}$, where $M$ and $N$ denote the number of test takers and questions, respectively. Each entry $Y_{ij}$ indicates the $i$-th test taker’s answer to the $j$-th question. Based on this response matrix, one can estimate the ability parameters $\theta$ and the difficulty parameters $d$ using a variety of statistical inference methods, such as the Bayesian inference via Markov chain Monte Carlo (MCMC) sampling (Chen et al., 2021) and variational inference (VI) (Blei et al., 2017). In this paper, we apply the VI method provided by py-irt (Lalor and Rodriguez, 2023). Concretely, we fit the Rasch model via stochastic variational inference on the response matrix with a mean-field Normal guide for $\theta$ and $d$, maximizing the evidence lower bound $\mathcal{L}$:
@@
$\max_{g}\ \mathcal{L}=$ $\sum_{i=1}^{M}\sum_{j=1}^{N}\mathbb{E}_{g}\!\left[Y_{ij}(\theta_{i}-d_{j})-\log\!\big(1+e^{\theta_{i}-d_{j}}\big)\right]$
$-\mathrm{KL}\!\big(g\,\|\,p\big)$
@@ 4 Method
In this section, we elaborate on the details of RIDE-Math. Inspired by the IRT-based difficulty estimation method with LLMs acting as students (Truong et al., 2025), we construct a difficulty ranker for math questions by performing parameter estimation described in Section 3. This ranker serves as a reward model by providing difficulty reward signals, which are used to guide reinforcement learning on the existing high-difficulty dataset DeepMath (He et al., 2025). Through this rewriting process, we are able to generate difficulty in evolving math reasoning data. The overall architecture of our method is shown in Figure 2.
Figure 2: The overall architecture of RIDE. RIDE trains a difficulty ranker based on IRT estimations to provide reward signals, guiding reinforcement learning to drive the LLM to rewrite original math questions into variations that expose model’s reasoning vulnerabilities.
@@ 4.1.1 IRT Parameter Estimation
We perform IRT estimation based on the response matrix $\mathbf{Y}$. Following the parameter estimation method described in Section 3, we obtain an IRT-estimated difficulty $d$ for each question and construct $N$ question-difficulty pairs $(q,d)$ for training the difficulty ranker, where $q$ denotes the question.
@@ 4.1.2 Pairwise Difficulty Ranker
In traditional IRT, estimating the difficulty of a new question requires collecting fresh student response data. To endow IRT difficulty estimation with generalization to unseen questions, we adopt an amortized approach (Truong et al., 2025), which extracts content features via an embedding model and learn a mapping from question content to difficulty value. However, due to the domain-specific symbolic semantics of math questions (Mirzadeh et al., 2025), directly regressing from embeddings to a scalar difficulty may result in unreliable performance. Therefore, instead of regressing the difficulty value, we frame the task as a pairwise ranking question, training a ranker to identify the more difficult question from a pair of question embeddings. We use OpenAI’s text-embedding-3-large  [^1] https://platform.openai.com/docs/models/text-embedding-3-large model to obtain embedding vectors for each question-difficulty pair $(q,d)$, creating $N$ question-embedding-difficulty triplets as $(q,e,d)$. We then group these pairs by fine-grained topics (Gao et al., 2025) like algebra, geometry, and so on. Within each subcategory, ranking data are generated by forming pairwise combinations of questions. Specifically, for any two questions $(q_{i},q_{j})$ within the same subcategory, we create the ordered pair $(e_{i},e_{j},r_{ij})$ and its symmetric counterpart $(e_{j},e_{i},1-r_{ij})$ to reduce positional bias, where $r_{ij}$ denotes to the ranking label: it equals to $1$ if the question $i$’s difficulty $d_{i}$ exceeds the question $j$’s difficulty $d_{j}$, and 0 otherwise. To account for the influence of the magnitude gap between difficulty scores $d$, we assign a weight to each ranking pair. For $(e_{i},e_{j},r_{ij})$, the weight is $w_{ij}=\sqrt{\lvert d_{i}-d_{j}\rvert}$. We use these pairwise samples to train a multilayer perceptron (MLP) as the difficulty ranker. The input $\mathbf{x}$ of MLP is a fusion of absolute and relative features, which is concatenated as:
@@
$\mathcal{L}_{\mathrm{BCE}}=$ $\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\mathcal{B}}w_{ij}\Big(-r_{ij}\log p(i,j)$
$-(1-r_{ij})\log\!\big(1-p(i,j)\big)\Big)$